Larry is twice as old as Jane Jane is 6years younger then Tony Tony is 12 years younger than larry how old is each
Answer by satyareddy22(84)

Let L be the age of larry,J be the age of Jane,T be the age of Tony.
L=2J----(given)
T-J=6--------(1)
L-T=12-------(2)
substituting L=2J in (2)
2J-T=12------(3)
Now solving (2) &(3)
We get,
J=18
L=36
T=24